# Enhancing Robustness in Deep Reinforcement Learning: A Lyapunov Exponent Approach

### v1

@@ Abstract
Deep reinforcement learning agents achieve state-of-the-art performance in a wide range of simulated control tasks. However, successful applications to real-world problems remain limited. One reason for this dichotomy is because the learned policies are not robust to observation noise or adversarial attacks. In this paper, we investigate the robustness of deep RL policies to a single small state perturbation in deterministic continuous control tasks. We demonstrate that RL policies can be deterministically chaotic as small perturbations to the system state have a large impact on subsequent state and reward trajectories. This unstable non-linear behaviour has two consequences: First, inaccuracies in sensor readings, or adversarial attacks, can cause significant performance degradation; Second, even policies that show robust performance in terms of rewards may have unpredictable behaviour in practice. These two facets of chaos in RL policies drastically restrict the application of deep RL to real-world problems. To address this issue, we propose an improvement on the successful Dreamer V3 architecture, implementing a Maximal Lyapunov Exponent regularisation. This new approach reduces the chaotic state dynamics, rendering the learnt policies more resilient to sensor noise or adversarial attacks and thereby improving the suitability of Deep Reinforcement Learning for real-world applications.
@@ 1 Introduction
Deep neural networks (DNNs) have revolutionised reinforcement learning (RL) [25], enabling agents to excel in a diverse set of simulated control tasks [12, 16, 22, 23, 24]. However, trained deep RL policies are not robust controllers as DNNs are vulnerable to adversarial attacks [7, 26]. Adding a small amount of noise to each observation can cause these policies to make poor decisions, considerably degrading their overall performance [10, 11, 13]. This lack of stability poses a significant threat when applying deep RL to real-world environments where inaccurate sensors can easily introduce noise [5]. In this work, we argue that even high-performing deep RL policies are not robust controllers as they can create a chaotic closed-loop control system [4, 14]. These systems are characterised by a high sensitivity to initial conditions with small changes in initial system states producing vastly different long-term outcomes.
@@
    Even for high-performing policies with stable returns, it remains impossible to accurately predict the long-term behaviour of these chaotic control systems as they rely on noisy partially observable sensors to attain an observation. This unpredictability means safe and reliable behaviour cannot be guaranteed.
@@
Figure 1: Reward attained when a trained deterministic Soft Actor-Critic [8] agent controls the deterministic Walker Walk environment. Each system has the same initial configuration other than the torso angle which is perturbed by ±$\pm 5\times 10^{-4}$ degrees. This small perturbation causes the systems to significantly diverge after 50 steps due to the chaotic nature of the control interaction. Consequently, this affects overall performance as there is significant variation in the total reward attained.
@@ 2.1 Robust reinforcement learning
$J(\theta)=\mathbb{E}_{s\sim\rho_{0}}\left[\sum_{t=0}^{\infty}\gamma^{t}\times r(s_{t},a_{t})\right]$ (1)
RL policies are said to be robust if they can maintain consistent behaviour and reliable performance in the face of noise or adversarial attacks. This stability is crucial for the safe deployment of deep RL to real-world applications where observation noise is inevitable. Recently, a wide range of attack methods have been developed which can easily compromise the performance of deep RL policies with relatively low levels of intervention. In their seminal work, Huang et al [10] extended the Fast Gradient Sign Method [7] attack to deep RL policies, demonstrating the addition of a small amount of strategic noise to every observation is sufficient to degrade the performance of trained deep RL agents. Furthermore, Kos & Song [11] showed this attack does not need to occur at every step as less frequent attacks with small Gaussian noise still produce a drop in performance. Additionally, they demonstrated that only perturbing the observation when the value function surpasses a set threshold still produces poor performance while significantly decreasing the frequency of the attack. A similar result was established by Lin et al [13] who showed that specifically attacking when the relative performance gain between best and worst action surpasses a defined threshold can successfully degrade performance. These findings demonstrate that consistent and inconsistent small perturbations to the system can easily confuse deep RL policies resulting in unintended and detrimental behaviour. Note that small observation noise is frequent in real-world systems, and therefore, this lack of robustness severely limits the application of deep RL to real-world environments.
Table 1: Stabilty of a dynamical system for different values of the Maximal Lyapunov Exponent ($\lambda_{1}$) and Sum of Lyapunov Exponents ($\lambda_{\Sigma}$).
@@ 2.2 Measuring stability: Lyapunov Exponents
Lyapunov Exponents (LEs) [15, 20] provide a method for quantifying the stability of complex, non-linear, high-dimensional systems by measuring the deformation rate of a small hyperellipsoid under the effects of a transition function. In general, for a dynamical system with $N$ degrees of freedom, there are $N$ LEs each representing the exponential growth rate of a unique principal axis of the hyperellipsoid. Given a set of $N$ ordered exponents ($\lambda_{1}\geq\lambda_{2}\geq...\geq\lambda_{N}$), the volume of the hyperellipsoid grows proportional to $e^{(\lambda_{1}+\lambda_{2}+...\lambda_{N})t}$ and the length grows proportionally to $e^{\lambda_{1}t}$. From this definition, the Maximal Lyapunov Exponent (MLE) (Equation 2) and the Sum of Lyapunov Exponents (SLE) (Equation 3) are used to determine if a system is stable, chaotic or unstable as outlined in Table 1.
$\lambda_{1}=\lim_{t\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{1}{t}\ln\left(\frac{|s_{t}~{}-~{}\hat{s}_{t}|}{\epsilon}\right)$ (2)
@@
Dynamical systems with a negative MLE ($\lambda_{1}\leq 0$) are stable as all principle axes of the hyperellipsoid exponentially decrease to zero [21]. In these systems, any trajectories produced from similar initial positions converge to the same trajectory given sufficient time. Conversely, systems with positive MLE ($\lambda_{1}>0$) and SLE ($\lambda_{\Sigma}>0$) are unstable as the resulting state trajectories diverge at an exponential rate. However, for a positive MLE ($\lambda_{1}>0$) and negative SLE ($\lambda_{\Sigma}<0$), similar trajectories will diverge at an exponential rate but remain confined to a subregion of the phase space known as a chaotic attractor [4, 14]. This bounded exponential divergence means trajectories in this region of the state space are unstable and only replicable given the exact same starting state: small perturbations to the starting state will produce significantly different long-term outcomes which appear random and uncorrelated. As a result, it is impossible to predict the long-term behaviour of a chaotic system given an approximation of the initial state.
@@
To measure the stability of a known dynamical system, the full spectrum of Lyapunov Exponents can be estimated using the approach outlined by Benettin et al [2, 3]. This method represents the spectrum as a set of small perturbation vectors which are iteratively updated using the known transition function. To avoid all vectors collapsing in the direction of maximal growth, they are periodically Gram-Schmidt orthonormalized so that each vector maintains a unique direction. Performing this orthonormalization allows for the detection of both positive and negative Lyapunov exponents up to the dimension of the phase space. The Spectrum of Lyapunov Exponents are then determined as the average log rate of divergence of the perturbation vectors as they are Gram-Schmidt orthonormalized. Estimation of the full spectrum of Lyapunov Exponents using this method allows for the estimation of $\lambda_{1}$ and $\lambda_{\Sigma}$ which can quantify the stability of the dynamical system.
2.3 Chaotic in reinforcement learning
@@
Previous studies have investigated the chaotic state and reward dynamics produced by Reinforcement Learning policies. Rahn et al [19] showed the policy optimisation landscape can contain high-frequency discontinuities in the vicinity of a trained policy. A similar result was established by Wang et al [28], who proved that control systems with Lipschitz continuous reward and transition functions only have a Lipschitz continuous objective function if $\lambda_{1}<-\ln(\gamma)$. When $\lambda_{1}>-\ln(\gamma)$, the objective function is a fractal and is $\alpha$-Hölder continuous with holder exponent $\alpha=-\ln(\gamma)/\lambda_{1}$. Consequently, a single update to the policy in these chaotic control systems can produce substantially different total rewards. Furthermore, Parmas et al [17] demonstrated that chaos exists in model-based RL methods due to repeated nonlinear predictions and this instability causes gradients to explode during training.
@@
While these works use chaos theory to highlight an important issue in the field of RL policy learning, they are focused primarily on the stability of the policy subject to policy parameter perturbations during training. Our work uses similar concepts but instead focuses on the stability of fully trained RL policies subject to state perturbations and the adverse effect this has on the total reward attained in realistic environments. To our knowledge, we are the first to use the Lyapunov Exponent to estimate the level of chaos produced by trained DNN policies in continuous control tasks.
@@
Figure 2: Total episode reward for the Pointmass Easy (PM), Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (SW), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) environments when controlled by trained instances of SAC, TD3, Dreamer V3 (DR3) and an agent which takes no actions (None). Each policy-environment combination is independently trained with three random seeds and the average interquartile episode reward with a bootstrapped 95% confidence interval is reported over 80 evaluation episodes each with a fixed length of 1000 steps.
@@ 3 Chaotic state dynamics
In this section, we use Lyapunov Exponents to identify the level of chaos produced by various state-of-the-art deep reinforcement learning policies in continuous control environments. We claim that the presence of chaos in the MDP implies that the policies are not robust controllers as trivial changes to the system state produce significantly different long-term state trajectories. This instability poses a significant problem for real-world control systems where consistent and predictable behaviour is necessary.
@@
To closely match real-world applications we estimate the stability of tasks from the DeepMind Control Suite [27] when controlled by deep RL policies. These simulated environments provide a range of deterministic control problems with continuous state spaces as outlined in Appendix A.1. For each control task we independently train three instances of Soft Actor-Critic (SAC) [8], Twin Delayed Deep Deterministic Policy Gradients (TD3) [6] and Dreamer V3 [9] as these represent state-of-the-art off-policy and model-based methods for continuous control tasks. Furthermore, to determine if trained deep RL policies directly influence the stability of each control system, we also introduce a passive controller that takes no actions. All models are based on the Stable Baselines 3 [18] implementation with the parameters outlined in Appendix A.2 and trained using an Intel Core i7-8700 CPU workstation with an Nvidia RTX 2080 Ti GPU and 32GB of RAM. The average interquartile reward and bootstrapped 95% confidence interval [1] for each policy type is reported in Figure 2. These results show that all three algorithms learn policies which performed well and significantly better than the no-action baseline. However, this does not speak to the types of dynamics produced or how robust these policies are to local state perturbations.
@@
Figure 3: Estimated Maximal Lyapunov Exponent (MLE) and Sum of Lyapunov Exponents (SLE) for the Pointmass (PM), Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (SW), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) environments when controlled by a trained instance of SAC, TD3, Dreamer V3 (DR3) and an agent which takes no actions (None). Each policy-environment combination is independently trained with three random seeds and the interquartile average MLE & SLE for each seed is calculated using 20 initial states. A bootstrapped 95% confidence interval is included to show the variation in MLE and SLE across random seeds.
@@
To identify the stability of each policy-environment interaction we estimate the full spectrum of Lyapunov Exponents using the method proposed by Benettin et al  [2, 3]. The spectrum is calculated using states sampled from the initial state distribution for each environment. A perturbation vector is initialised for each state dimension at a distance of $10^{-4}$ from the sample state. This represents an arbitrarily small change to the control system without introducing numerical precision errors. The spectrum is calculated over 1 thousand environment steps and perturbation vectors are otheronormalised every ten steps to prevent divergence saturation. This process is repeated with 20 initial states and the average value for each exponent is used to calculate the MLE and SLE. Ablation studies for these constants are provided in Appendix B.
@@
Figure 3 provides the bootstrapped 95% confidence interval for the MLE and SLE produced by each policy-environment pair. These results indicate that all the environments are naturally invariant to small perturbations as the No Action baseline has $\lambda_{1}=0$. Conversely, when these environments are controlled by deep RL policies, $\lambda_{1}$ can be non-zero indicating DNN controllers directly influence the stability of these control systems.
@@
We find SAC and TD3 produce stable state dynamics in simple low-dimensional environments (Pointmass, Cartpole Balance and Cartpole Swingup) as they have negative MLE. Therefore, if a small perturbation is made to any of the states in these environments, the subsequent state trajectories converge. This result is consistent with the definition of the reward function for each of these tasks as they provide high rewards for maintaining the system at a fixed location. However, when controlled by Dreamer V3, these simple systems exhibit low levels of chaos as they have positive MLE and negative SLE. As a result, state trajectories in these environments are highly sensitive to initial conditions with similar states producing significantly different long-term outcomes as shown in Figure 4(a). Instead of converging to a single state, these trajectories exhibit chaotic behaviour, continuously orbiting within a region which yields high rewards.
@@
Furthermore, all deep RL methods produce chaotic dynamics in the complex high-dimensional environments (Walker Stand, Walker Walk, Walker Run and Cheetah Run) as indicated by the positive MLE and negative SLE. This means these policies cannot account for arbitrarily small changes in the system’s state since any small changes produce exponentially diverging state trajectories, as illustrated in Figure 4(b). This poses a significant problem for real-world applications of RL as observation perturbations are easily introduced via imperfect measurements or sensor noise. Therefore, for these complex high-dimensional environments controlled by RL policies, it is impossible to guarantee stability as the system cannot correct itself after even a single small observation inaccuracy.

### v2

@@ Abstract
Deep reinforcement learning agents achieve state-of-the-art performance in a wide range of simulated control tasks. However, successful applications to real-world problems remain limited. One reason for this dichotomy is because the learnt policies are not robust to observation noise or adversarial attacks. In this paper, we investigate the robustness of deep RL policies to a single small state perturbation in deterministic continuous control tasks. We demonstrate that RL policies can be deterministically chaotic, as small perturbations to the system state have a large impact on subsequent state and reward trajectories. This unstable non-linear behaviour has two consequences: first, inaccuracies in sensor readings, or adversarial attacks, can cause significant performance degradation; second, even policies that show robust performance in terms of rewards may have unpredictable behaviour in practice. These two facets of chaos in RL policies drastically restrict the application of deep RL to real-world problems. To address this issue, we propose an improvement on the successful Dreamer V3 architecture, implementing Maximal Lyapunov Exponent regularisation. This new approach reduces the chaotic state dynamics, rendering the learnt policies more resilient to sensor noise or adversarial attacks and thereby improving the suitability of deep reinforcement learning for real-world applications.
@@ 1 Introduction
Deep neural networks (DNNs) have revolutionised reinforcement learning (RL) [25], enabling agents to excel in a diverse set of simulated control tasks [12, 16, 22, 23, 24]. However, trained deep RL policies are not robust controllers as DNNs are vulnerable to adversarial attacks [7, 26]. Adding a small amount of noise to each observation can cause these policies to make poor decisions, considerably degrading their overall performance [10, 11, 13]. This lack of stability poses a significant threat when applying deep RL to real-world environments, where inaccurate sensors can easily introduce noise [5]. In this work, we argue that even high-performing deep RL policies are not robust controllers as they can create a chaotic closed-loop control system [4, 14]. These systems are characterised by a high sensitivity to initial conditions, with small changes in initial system states producing vastly different long-term outcomes.
@@
    Even for high-performing policies with stable returns, it remains impossible to accurately predict the long-term behaviour of these chaotic control systems as they rely on noisy partially observable sensors to attain an observation. This unpredictability means that safe and reliable behaviour cannot be guaranteed.
@@
Figure 1: Reward attained when a trained deterministic Soft Actor-Critic [8] agent controls the deterministic Walker Walk environment. Each system has the same initial configuration other than the torso angle, which is perturbed by ±$\pm 5\times 10^{-4}$ degrees. This small perturbation causes the systems to significantly diverge after 50 steps due to the chaotic nature of the control interaction. Consequently, this affects overall performance as there is significant variation in the total reward attained.
@@ 2.1 Robust reinforcement learning
$J(\theta)=\mathop{\mathbb{E}}_{s_{0}\sim\rho_{0}}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}\times r(s_{t},a_{t})~{}\right|~{}{s_{t+1}=f(s_{t},a_{t}),~{}a_{t}=\pi_{\theta}(s_{t})}\right]$ (1)
RL policies are said to be robust if they can maintain consistent behaviour and reliable performance in the face of noise or adversarial attacks. This stability is crucial for the safe deployment of deep RL to real-world applications where observation noise is inevitable. Recently, a wide range of attack methods have been developed which can easily compromise the performance of deep RL policies with relatively low levels of intervention. In their seminal work, Huang et al. [10] extended the Fast Gradient Sign Method [7] attack to deep RL policies, demonstrating that the addition of a small amount of strategic noise to every observation is sufficient to degrade the performance of trained deep RL agents. Furthermore, Kos & Song [11] showed that this attack does not need to occur at every step, as less frequent attacks with small Gaussian noise still produce a drop in performance. Additionally, they demonstrated that only perturbing the observation when the value function surpasses a set threshold still produces poor performance while significantly decreasing the frequency of the attack. A similar result was established by Lin et al. [13], who showed that specifically attacking when the relative performance gain between best and worst action surpasses a defined threshold can successfully degrade performance. These findings demonstrate that consistent and inconsistent small perturbations to the system can easily confuse deep RL policies, resulting in unintended and detrimental behaviour. Note that small observation noise is frequent in real-world systems, and therefore, this lack of robustness severely limits the application of deep RL to real-world environments.
Table 1: Stability of a dynamical system for different values of the Maximal Lyapunov Exponent ($\lambda_{1}$) and Sum of Lyapunov Exponents ($\lambda_{\Sigma}$).
@@ 2.2 Measuring stability: Lyapunov Exponents
Lyapunov Exponents [15, 20] provide a method for quantifying the stability of complex, non-linear, high-dimensional systems by measuring the deformation rate of a small hyperellipsoid under the effects of a transition function. In general, for a dynamical system with $N$ degrees of freedom, there are $N$ Lyapunov Exponents, each representing the exponential growth rate of a unique principal axis of the hyperellipsoid. Given a set of $N$ ordered exponents ($\lambda_{1}\geq\lambda_{2}\geq...\geq\lambda_{N}$), the volume of the hyperellipsoid grows proportionally to $e^{(\lambda_{1}+\lambda_{2}+...\lambda_{N})t}$ and the length grows proportionally to $e^{\lambda_{1}t}$. From this definition, the Maximal Lyapunov Exponent (MLE) (Equation 2) and the Sum of Lyapunov Exponents (SLE) (Equation 3) are used to determine if a system is stable, chaotic or unstable, as outlined in Table 1.
$\lambda_{1}=\lim_{t\rightarrow\infty}~{}\lim_{\hat{s}_{0}\rightarrow s_{0}}~{}\frac{1}{t}\ln\left(\frac{|s_{t}~{}-~{}\hat{s}_{t}|}{|s_{0}~{}-~{}\hat{s}_{0}|}\right)$ (2)
@@
Dynamical systems with a negative MLE ($\lambda_{1}\leq 0$) are stable as all principal axes of the hyperellipsoid exponentially decrease to zero [21]. In these systems, any trajectories produced from similar initial positions converge to the same trajectory given sufficient time. Conversely, systems with positive MLE ($\lambda_{1}>0$) and SLE ($\lambda_{\Sigma}>0$) are unstable as the resulting state trajectories diverge at an exponential rate. However, for a positive MLE ($\lambda_{1}>0$) and negative SLE ($\lambda_{\Sigma}<0$), similar trajectories will diverge at an exponential rate but remain confined to a subregion of the phase space known as a chaotic attractor [4, 14]. This bounded exponential divergence means trajectories in this region of the state space are unstable and only replicable given the exact same starting state: small perturbations to the starting state will produce significantly different long-term outcomes which appear random and uncorrelated. As a result, it is impossible to predict the long-term behaviour of a chaotic system given an approximation of the initial state.
@@
To measure the stability of a known dynamical system, the full spectrum of Lyapunov Exponents can be estimated using the approach outlined by Benettin et al. [2, 3]. This method represents the spectrum as a set of small perturbation vectors which are iteratively updated using the known transition function. To avoid all vectors collapsing in the direction of maximal growth, they are periodically Gram-Schmidt orthonormalised so that each vector maintains a unique direction. Performing this orthonormalisation allows for the detection of both positive and negative Lyapunov exponents up to the dimension of the phase space. The spectrum of Lyapunov Exponents is then determined as the average log rate of divergence of the perturbation vectors as they are Gram-Schmidt orthonormalised. Estimation of the full spectrum of Lyapunov Exponents using this method allows for the estimation of $\lambda_{1}$ and $\lambda_{\Sigma}$ which can quantify the stability of the dynamical system.
2.3 Chaos in reinforcement learning
@@
Previous studies have investigated the chaotic state and reward dynamics produced by reinforcement learning policies. Rahn et al. [19] showed that the policy optimisation landscape can contain high-frequency discontinuities in the vicinity of a trained policy. A similar result was established by Wang et al. [28], who proved that control systems with Lipschitz continuous reward and transition functions only have a Lipschitz continuous objective function if $\lambda_{1}<-\ln(\gamma)$. When $\lambda_{1}>-\ln(\gamma)$, the objective function is a fractal and is $\alpha$-Hölder continuous with holder exponent $\alpha=-\ln(\gamma)/\lambda_{1}$. Consequently, a single update to the policy in these chaotic control systems can produce substantially different total rewards. Furthermore, Parmas et al. [17] demonstrated that chaos exists in model-based RL methods due to repeated nonlinear predictions and this instability causes gradients to explode during training.
@@
While these works use chaos theory to highlight an important issue in the field of RL policy learning, they are focused primarily on the stability of the policy subject to policy parameter perturbations during training. Our work uses similar concepts but instead focuses on the stability of fully trained RL policies subject to state perturbations and the adverse effect this has on the total reward attained in realistic environments. To our knowledge, we are the first to use the spectrum of Lyapunov Exponents to estimate the level of chaos produced by trained DNN policies in continuous control tasks.
@@
Figure 2: Total episode reward for the Pointmass Easy (PM), Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (WS), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) environments when controlled by trained instances of SAC, TD3, Dreamer V3 (DR3) and an agent which takes no actions (None). Each policy-environment combination is independently trained with three random seeds and the average interquartile episode reward with a bootstrapped 95% confidence interval is reported over 80 evaluation episodes each with a fixed length of 1000 steps.
@@ 3 Chaotic state dynamics
In this section, we use Lyapunov Exponents to identify the level of chaos produced by various state-of-the-art deep reinforcement learning policies in continuous control environments. We claim that the presence of chaos in the MDP implies that the policies are not robust controllers, as trivial changes to the system state produce significantly different long-term state trajectories. This instability poses a significant problem for real-world control systems where consistent and predictable behaviour is necessary.
@@
To closely match real-world applications, we estimate the stability of tasks from the DeepMind Control Suite [27] when controlled by deep RL policies. These simulated environments provide a range of deterministic control problems with continuous state spaces, as outlined in Appendix A.1. For each control task we independently train three instances of Soft Actor-Critic (SAC) [8], Twin Delayed Deep Deterministic Policy Gradients (TD3) [6] and Dreamer V3 [9], as these represent state-of-the-art off-policy and model-based methods for continuous control tasks. Furthermore, to determine if trained deep RL policies directly influence the stability of each control system, we also introduce a passive controller that takes no actions. All models are based on the Stable Baselines 3 [18] implementation with the parameters outlined in Appendix A.2 and trained using an Intel Core i7-8700 CPU workstation with an Nvidia RTX 2080 Ti GPU and 32GB of RAM. The average interquartile reward and bootstrapped 95% confidence interval [1] for each policy type are reported in Figure 2. These results show that all three algorithms learn policies which performed well and significantly better than the no-action baseline. However, this does not speak to the types of dynamics produced or how robust these policies are to local state perturbations.
@@
Figure 3: Estimated Maximal Lyapunov Exponent (MLE) and Sum of Lyapunov Exponents (SLE) for the Pointmass (PM), Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (WS), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) environments when controlled by a trained instance of SAC, TD3, Dreamer V3 (DR3) and an agent which takes no actions (None). Each policy-environment combination is independently trained with three random seeds and the interquartile average MLE & SLE for each seed is calculated using 20 initial states. A bootstrapped 95% confidence interval is included to show the variation in MLE and SLE across random seeds.
@@
To identify the stability of each policy-environment interaction, we estimate the full spectrum of Lyapunov Exponents using the method proposed by Benettin et al.  [2, 3]. The spectrum is calculated using states sampled from the initial state distribution for each environment. A perturbation vector is initialised for each state dimension at a distance of $10^{-4}$ from the sample state. This represents an arbitrarily small change to the control system without introducing numerical precision errors. The spectrum is calculated over 1000 environment steps and perturbation vectors are orthonormalised every 10 steps to prevent divergence saturation. This process is repeated with 20 initial states and the average value for each exponent is used to calculate the MLE and SLE. Ablation studies for these constants are provided in Appendix B.
@@
Figure 3 provides the bootstrapped 95% confidence interval for the MLE and SLE produced by each policy-environment pair. These results indicate that all the environments are naturally invariant to small perturbations as the no action baseline has $\lambda_{1}=0$. Conversely, when these environments are controlled by deep RL policies, $\lambda_{1}$ can be non-zero, indicating that DNN controllers directly influence the stability of these control systems.
@@
We find that SAC and TD3 produce stable state dynamics in simple low-dimensional environments (Pointmass, Cartpole Balance and Cartpole Swingup) as they have negative MLE. Therefore, if a small perturbation is made to any of the states in these environments, the subsequent state trajectories converge. This result is consistent with the definition of the reward function for each of these tasks as they provide high rewards for maintaining the system at a fixed location. However, when controlled by Dreamer V3, these simple systems exhibit low levels of chaos as they have positive MLE and negative SLE. As a result, state trajectories in these environments are highly sensitive to initial conditions, with similar states producing significantly different long-term outcomes as shown in Figure 4(a). Instead of converging to a single state, these trajectories exhibit chaotic behaviour, continuously orbiting within a region which yields high rewards.
@@
Furthermore, all deep RL methods produce chaotic dynamics in the complex high-dimensional environments (Walker Stand, Walker Walk, Walker Run and Cheetah Run), as indicated by the positive MLE and negative SLE. This means that these policies cannot account for arbitrarily small changes in the system’s state since small changes produce exponentially diverging state trajectories, as illustrated in Figure 4(b). This poses a significant problem for real-world applications of RL, as observation perturbations are easily introduced via imperfect measurements or sensor noise. Therefore, for these complex high-dimensional environments controlled by deep RL policies, it is impossible to guarantee stability as the system cannot correct itself after a single inaccurate observation.